# Continuous-time q-learning in jump-diffusion models under Tsallis entropy

### v3

@@ Abstract
This paper studies the continuous-time reinforcement learning in jump-diffusion models by featuring the q-learning (the continuous-time counterpart of Q-learning) under Tsallis entropy regularization. Contrary to the Shannon entropy, the general form of Tsallis entropy renders the optimal policy not necessarily a Gibbs measure. Herein, the Lagrange multiplier and KKT condition are needed to ensure that the learned policy is a probability density function. As a consequence, the characterization of the optimal policy using the q-function also involves a Lagrange multiplier. In response, we establish the martingale characterization of the q-function and devise two q-learning algorithms depending on whether the Lagrange multiplier can be derived explicitly or not. In the latter case, we consider different parameterizations of the optimal q-function and the optimal policy, and update them alternatively in an Actor-Critic manner. We also study two numerical examples, namely, an optimal liquidation problem in dark pools and a non-LQ control problem. It is interesting to see therein that the optimal policies under the Tsallis entropy regularization can be characterized explicitly, which are distributions concentrated on some compact support. The satisfactory performance of our q-learning algorithms is illustrated in each example.
@@ 1 Introduction
Reinforcement learning (RL) has witnessed fast-growing advancements in recent years, especially in the continuous-time framework. Q-learning algorithm (see Watkins 1989, Watkins and Dayan 1992) is widely known as popular value-based learning method in the discrete-time framework. By learning a Q-function that maps state-action pairs to expected rewards, Q-learning enables agents to optimize policies by iteratively updating Q-values to reflect the expected returns. (Sutton 2018). However, learning the Q-function in a continuous-time setting is not straightforward, as the Q-function collapses to a value function independent of actions. Continuous-time models have been popularized in quantitative finance thanks to their merits that can effectively capture real-time adjustments of dynamics to the fast-changing environment and facilitate the characterization of more precise control. This is particularly advantageous in tasks requiring fine-grained decision making, such as high-frequency trading. Given that real-world decision-making often unfolds in continuous time, there has been growing interest and debate among researchers in developing effective continuous-time RL algorithms. Early research on continuous-time RL dates back to Baird (1994), who introduced ”advantage updating” for continuous-time control, and Munos and Bourgine (1997), who explored RL for continuous stochastic control problems. Recent pioneer studies Wang et al. (2020), Jia and Zhou (2022a, b, 2023) have laid the theoretical foundations for continuous-time RL with continuous state and action spaces. In particular, Jia and Zhou (2023) propose a continuous-time q-learning approach, generalizing the conventional Q-function and Q-learning algorithm from a discrete-time setting to continuous-time counterparts by utilizing the first-order approximation of the advantage function (the difference between the Q-function and the value function) with respect to time.
@@
In contrast, the present paper aims to develop a continuous-time q-learning method for jump-diffusion models under Tsallis entropy. Tsallis (1988) proposes a generalization of Shannon entropy that provides greater flexibility and robustness to handle learning tasks with diverse policy distributions especially for the purpose of concentrated sample actions. Particularly, Tsallis entropy is superior in scenarios with prevalent non-Gaussian, heavy-tailed behavior, on the compact support. As a direct consequence, the sampled actions are more concentrated in certain regions such that some extreme and risky decisions can be avoided during the learning procedure (see Mertikopoulos and Sandholm 2016, Chow et al. 2018). By adjusting its index parameter, Tsallis entropy regularization can turn the learned optimal policy into different types, offering greater flexibility in managing uncertainty and incentivizing exploration in RL. Lee et al. (2018, 2019) study a class of Markov decision processes (MDP) with Tsallis entropy maximization. Donnelly and Jaimungal (2024) recently investigate the optimal control in models with latent factors where the agent controls the distribution over actions by rewarding exploration with Tsallis entropy in both discrete and continuous time.
@@
Our paper then applies the proposed q-learning algorithms under Tsallis entropy regularization to two financial applications. The first example employs the q-learning method to solve an LQ control problem with pure jumps that optimizes the trading strategies in dark pools as in Kratz and Schöneborn (2014, 2015). When trading occurs concurrently in both the primary market and dark pools, the distribution of trades in these venues follows a two-dimensional random vector. Notably, the optimal policy in this LQ framework can be obtained explicitly, which is a non-Gaussian distribution with a compact support. The second examples adopts the q-learning method to solve a class of non-LQ jump-diffusion control problems related to selecting different repo rates (Bichuch et al. 2018). When dealing with two distinct repo rates, the trading proportions of these financial products are governed by a two-dimensional random vector. An interesting finding is that the optimal policy under a general power utility can be explicitly derived when the Tsallis entropy index equals $2$, but no explicit characterization of the optimal policy can be obtained under the conventional Shannon entropy, illustrating one technical advantage of Tsallis entropy over the Shannon entropy.
@@ 2.1 Exploratory formulation in reinforcement learning
For a fixed time horizon $T\in(0,\infty)$, let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered probability space with the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ satisfying the usual conditions. On this probability space, the process $W=(W_{t})_{t\in[0,T]}$ is a scalar Brownian motion and the process ${N}=(N(t,z);z\in\mathds{R})_{t\in[0,T]}$ is an $\mathbb{F}$-adapted Poisson point process with an intensity $\sigma$-finite measure $\nu$ on ${\cal B}(\mathds{R}\setminus\{0\})$ satisfying $\int_{\mathds{R}\setminus\{0\}}\min\{z^{2},1\}\nu(dz)<\infty$, which is independent of $W$. Let us consider the following controlled jump-diffusion process that, for $t\in(0,T]$,
$dX_{t}^{u}=b(t,X_{t}^{u},u_{t})dt+\sigma(t,X_{t}^{u},u_{t})dW_{t}+\int_{\mathds{R}}\varphi(t,X_{t-}^{u},u_{t},z)N(dt,dz),\quad X_{0}^{u}=x\in\mathds{R},$ (2.1)
where $u=(u_{t})_{t\in[0,T]}$ is an $\mathbb{F}$-predictable process taking values on $U\subset\mathds{R}^{d}$, and the set of admissible controls is denoted by $\mathcal{U}$. Here, $b(t,x,u):[0,T]\times\mathds{R}\times U\to\mathds{R}$, $\sigma(t,x,u):[0,T]\times\mathds{R}\times U\to\mathds{R}$ and $\varphi(t,x,u,z):[0,T]\times\mathds{R}\times U\times\mathds{R}\to\mathds{R}$ are assumed to be measurable functions.
@@
where $f(t,x,u):[0,T]\times\mathds{R}\times U\to\mathds{R}$ stands for the running reward function, $g(x):\mathds{R}\to\mathds{R}$ is the terminal reward function, and the expectation operator $\mathbb{E}_{t}^{\mathbb{P}}[\cdot]:=\mathbb{E}^{\mathbb{P}}\left[\cdot|X_{t}=x\right]$.
@@
Given the full knowledge of the coefficients $b,\sigma,\varphi,f,g$, and the intensity measure $\nu$ in (2.1)-(2.2), the classical methods such as the dynamic programming principle and the stochastic maximum principle can be employed to solve the above optimal control problem (2.1)-(2.2). However, in reality, the decision maker may have limited or no information about the environment (i.e., $b,\sigma,\varphi,f,g,\nu$ are unknown). The reinforcement learning approach provides an efficient way to learn the optimal control in (2.1)-(2.2) in the unknown environment through the repeated trial-and-error procedure by taking actions and interacting with the environment. Specifically, he tries a sequence of actions $u=(u_{t})_{t\in[0,T]}$ and observe the corresponding state process $X=(X^{u}_{t})_{t\in[0,T]}$ along with a stream of running rewards $(f(t,X^{u}_{t},u_{t}))_{t\in[0,T]}$ and the terminal reward $g(X_{T}^{u})$, and continuously update and improve his or her actions based on these observations.
@@
To describe the exploration step in reinforcement learning, we can randomize the action $u$ and consider its distribution. Let ${\cal P}(U)$ be the set of probability measures on $U$ and $\pi=(\pi_{t})_{t\in[0,T]}$ be a given stochastic feedback policy with $\pi_{t}=\pi(\cdot|t,x)\in{\cal P}(U)$ for any $t\in[0,T]$. For a Borel space $(E,{\cal B}(E))$, consider another probability space $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$ and a measurable function $\phi:[0,T]\times\mathbb{R}\times E\to U$ such that, for all $(t,x)\in[0,T]\times\mathbb{R}$, the $U$-valued random variable $\omega\to\phi(t,x,\xi(\omega))$ has the distribution $\pi(\cdot|t,x)$. Let $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$ and $(\Omega^{\xi_{n}},\mathcal{F}^{\xi_{n}},\mathbb{P}^{\xi_{n}},\xi_{n})_{n\in\mathbb{N}_{0}}$ be independent copies of $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi},\xi)$. Introduce the following probability space with the form given by
@@
$(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{Q}):=\left(\Omega\times\prod_{n=0}^{\infty}\Omega^{\xi_{n}},\mathcal{F}\otimes\bigotimes_{n=0}^{\infty}\mathcal{F}^{\xi_{n}},\mathbb{P}\otimes\bigotimes_{n=0}^{\infty}\mathbb{P}^{\xi_{n}}\right),$ (2.3)
@@
and denotes $\mathbb{F}^{\prime}=(\mathcal{F}_{t}\vee\mathcal{F}^{\prime}_{t})_{t\in[0,T]}$. Furthermore, consider the time grid $G_{t:T}:=\{t=s_{0}<s_{1}<\cdots<s_{n}=T\}$ and sample actions from $\pi$ only at the grid points in $G_{t:T}$. The state process satisfies, for all $i=0,\ldots,n-1$ and $s\in[s_{i},s_{i+1}]$,
$X_{s}$ $=X_{s_{i}}+\int_{s_{i}}^{s}b(\ell,X_{\ell},u_{s_{i}})d\ell+\int_{s_{i}}^{s}\sigma(\ell,X_{\ell},u_{s_{i}})dW_{\ell}+\int_{s_{i}}^{s}\int_{\mathds{R}}\varphi(\ell,X_{\ell-}^{u},u_{s_{i}},z)N(d\ell,dz),$
$u_{s_{i}}$ $=\phi(s_{i},X_{s_{i}},\xi_{i}).$ (2.4)
@@
$dX_{s}^{\pi}$ $=b(s,X_{s}^{\pi},u_{s}^{\pi})ds+\sigma(s,X_{s}^{\pi},u_{s}^{\pi})dW_{s}+\int_{\mathds{R}}\varphi(s,X_{s-}^{\pi},u_{s}^{\pi},z)N(ds,dz),$
$u_{s}^{\pi}$ $=u_{\delta(s)}\leavevmode\nobreak\ \text{with\leavevmode\nobreak\ $\delta(s)=s_{i}$ when $s\in[s_{i},s_{i+1})$}.$ (2.5)
Inspired by Wang et al. (2020), in which the Shannon entropy regularizer is introduced to encourage the exploration in RL, we here consider the so-called Tsallis entropy with order $p\geq 1$ as the regularizer for the same reason of policy exploration. Then, consider the following objective functional that
@@
$l_{p}(z)=\begin{cases}\displaystyle\frac{1}{p-1}(1-z^{p-1}),&p>1,\\[6.00006pt] \displaystyle\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\ln z,&p=1.\end{cases}$ (2.7)
@@
However, the representation (2.1)-(2.6) cannot be applied to derive exploratory HJB equation directly from the point of view of DPP. To this purpose, it is necessary to provide the relaxed version of the control problem. Inspired by Kushner (2000), let us introduce the process $\tilde{X}=(\tilde{X}_{s})_{s\in[t,T]}$ satisfying the following SDE, $\tilde{X}_{t}^{\pi}=x$, and for $s\in(t,T]$,
@@
$\quad+\int_{U}\int_{\mathds{R}}\varphi(s,\tilde{X}_{s-}^{\pi},u,z)\mathcal{N}(ds,dz,du),$ (2.8)
where $B=(B_{t})_{t\in[0,T]}$ is a scalar Brownian motion and $\mathcal{N}(ds,dz,du)$ is a Poisson random measure on $\mathds{R}_{+}\times{\mathds{R}}\times U$ with compensator $\nu(dz)\pi_{s}(u)duds$ independent of $B$. An interesting finding is that, for the pure jump controlled state model, the representation (2.1) of the relaxed controlled state process can be applied to derive exploratory HJB equations directly from the point of view of DPP, which is different from the controlled diffusion case as in Wang et al. (2020) in which the relaxed control form should be rewritten as an average formulation (in fact, for the diffusive case, the equivalence between the relaxed form and the average form). Therefore, we can formulate our reinforcement learning problem for the jump-diffusion controlled model (2.1) based on the relaxed control form (2.1). Thus, our reinforcement learning problem associated with the jump-diffusion controlled state process (2.1) can be stated as follows:
@@
${\rm s.t.}\leavevmode\nobreak\ \tilde{X}_{s}^{\pi}=x+\int_{t}^{s}\tilde{b}(\ell,\tilde{X}_{\ell}^{\pi},\pi_{\ell})d\ell+\int_{t}^{s}\tilde{\sigma}(\ell,\tilde{X}_{\ell}^{\pi},\pi_{\ell})dB_{\ell}+\int_{t}^{s}\int_{U}\int_{\mathds{R}}\varphi(\ell,\tilde{X}_{\ell-}^{\pi},u,z)\mathcal{N}(d\ell,dz,du).$
Here, $\Pi_{t}$ is the set of admissible (randomized) policies on $U$ and the coefficients $\tilde{b},\tilde{\sigma},\tilde{f},\tilde{l}_{p}$ are defined by, for $(s,x,\pi)\in[0,T]\times\mathds{R}\times{\cal P}(U)$,
$\tilde{b}(s,x,\pi):=\int_{U}b(s,x,u)\pi(u|s,x)du,\quad\tilde{\sigma}(s,x,\pi):=\sqrt{\int_{U}\sigma^{2}(s,x,u)\pi(u|s,x)du},$
$\tilde{f}(s,x,\pi):=\int_{U}f(s,x,u)\pi(u|s,x)du,\quad\tilde{l}_{p}(s,x,\pi):=\int_{U}l_{p}(\pi(u|s,x))\pi(u|s,x)du.$
@@
    $\pi(\cdot|s,x)\in{\cal P}(U)$ for all $(s,x)\in[t,T]\times\mathds{R}$, and $\pi(\cdot|\cdot,\cdot):U\times[t,T]\times\mathds{R}\to\mathds{R}$ is a measurable function;
@@
    $\int_{U}\left|\pi(u|s,x_{1})-\pi\left(u|s,x_{2}\right)\right|du\leq C\left|x_{1}-x_{2}\right|,\quad\forall x_{1},x_{2}\in\mathds{R}.$

### v4

@@ Abstract
This paper studies continuous-time reinforcement learning in jump-diffusion models by featuring the q-learning (the continuous-time counterpart of Q-learning) under Tsallis entropy regularization. Contrary to the Shannon entropy, the general form of Tsallis entropy renders the optimal policy not necessarily a Gibbs measure. Herein, the Lagrange multiplier and KKT condition are needed to ensure that the learned policy is a probability density function. As a consequence, the characterization of the optimal policy using the q-function also involves a Lagrange multiplier. In response, we establish the martingale characterization of the q-function and devise two q-learning algorithms depending on whether the Lagrange multiplier can be derived explicitly or not. We also study two numerical examples, namely, an optimal liquidation problem in dark pools and a non-LQ control problem. It is interesting to see therein that the optimal policies under the Tsallis entropy regularization can be characterized explicitly, which are distributions concentrated on some compact support. The satisfactory performance of our q-learning algorithms is illustrated in each example.
@@ 1 Introduction
Reinforcement learning (RL) has witnessed fast-growing advancements in recent years, especially in the continuous-time framework. Q-learning algorithm (see Watkins 1989, Watkins and Dayan 1992) is widely known as a popular value-based learning method in the discrete-time framework. By learning a Q-function that maps state-action pairs to expected rewards, Q-learning enables agents to optimize policies by iteratively updating Q-values to reflect the expected returns. (Sutton 2018). However, learning the Q-function in a continuous-time setting is not straightforward, as the Q-function collapses to a value function independent of actions. Continuous-time models have been popularized in quantitative finance thanks to their merits that can effectively capture real-time adjustments of dynamics to the fast-changing environment and facilitate the characterization of more precise control. This is particularly advantageous in tasks requiring fine-grained decision-making, such as high-frequency trading. Given that real-world decision-making often unfolds in continuous time, there has been growing interest and debate among researchers in developing effective continuous-time RL algorithms. Early research on continuous-time RL dates back to Baird (1994), who introduced ”advantage updating” for continuous-time control, and Munos and Bourgine (1997), who explored RL for continuous stochastic control problems. Recent pioneer studies Wang et al. (2020), Jia and Zhou (2022a, b, 2023) have laid the theoretical foundations for continuous-time RL with continuous state and action spaces. In particular, Jia and Zhou (2023) propose a continuous-time q-learning approach, generalizing the conventional Q-function and Q-learning algorithm from a discrete-time setting to continuous-time counterparts by utilizing the first-order approximation of the advantage function (the difference between the Q-function and the value function) with respect to time.
@@
In contrast, the present paper aims to develop a continuous-time q-learning method for jump-diffusion models with state-dependent intensity under Tsallis entropy. Tsallis (1988) proposes a generalization of Shannon entropy that provides greater flexibility and robustness to handle learning tasks with diverse policy distributions especially for the purpose of concentrated sample actions. Particularly, Tsallis entropy is superior in scenarios with prevalent non-Gaussian, heavy-tailed behavior, on the compact support. As a direct consequence, the sampled actions are more concentrated in certain regions such that some extreme and risky decisions can be avoided during the learning procedure (see Mertikopoulos and Sandholm 2016, Chow et al. 2018). By adjusting its index parameter, Tsallis entropy regularization can turn the learned optimal policy into different types, offering greater flexibility in managing uncertainty and incentivizing exploration in RL. Lee et al. (2018, 2019) study a class of Markov decision processes (MDP) with Tsallis entropy maximization. Donnelly and Jaimungal (2024) recently investigate the optimal control in models with latent factors where the agent controls the distribution over actions by rewarding exploration with Tsallis entropy in both discrete and continuous time.
@@
Our paper then applies the proposed q-learning algorithms under Tsallis entropy regularization to two financial applications. The first example employs the q-learning method to solve an LQ control problem with pure jumps that optimizes the trading strategies in dark pools as in Kratz and Schöneborn (2014, 2015). When trading occurs concurrently in both the primary market and dark pools, the distribution of trades in these venues follows a two-dimensional random vector. Notably, the optimal policy in this LQ framework can be obtained explicitly, which is a non-Gaussian distribution with a compact support. The second example adopts the q-learning to solve a class of non-LQ jump-diffusion control problems related to selecting different repo rates (Bichuch et al. 2018). When dealing with two distinct repo rates, the trading proportions of these financial products are governed by a two-dimensional random vector. An interesting finding is that the optimal policy under a general power utility can be explicitly derived when the Tsallis entropy index is set to be $2$, but no explicit characterization of the optimal policy can be obtained under the conventional Shannon entropy, illustrating one technical advantage of Tsallis entropy over the Shannon entropy. However, we emphasize that the goal of this study on Tsallis entropy is not to demonstrate better learning performance over Shannon entropy in convergence or accuracy. Rather, the motivation is to understand how Tsallis entropy may help to derive some optimal policies on compact support, which can be appealing in practical applications thanks to the sampled actions in bounded domain without ad-hoc truncations.
@@
Notations. Throughout this paper, we will use the following notations frequently. For $a,b\in\mathds{R}^{n}$, let $a\cdot b:=a^{\top}b$. For a matrix $A\in\mathbb{R}^{n\times d}$, we use $A^{\top}$ for its transpose, $\operatorname{tr}(A)$ for its trace, $|A|$ for its Euclidean (Frobenius) norm, and we write $A^{2}:=AA^{\top}$. For $A,B\in\mathbb{R}^{n\times d}$, let $A\circ B:=\operatorname{tr}(AB^{\top})$. If $A$ is positive semidefinite, its square root $\sqrt{A}$ is defined via the singular value decomposition $A=UBV^{\top}$, where $U,V$ are orthogonal and $B$ is diagonal. We set $\sqrt{A}=UB^{1/2}V^{\top}$, with $B^{1/2}$ denoting the diagonal matrix of square roots of the entries of $B$. For an open set $\mathcal{O}\subset\mathbb{R}^{n}$, $C^{k}(\mathcal{O})$ is the space of real functions on $\mathcal{O}$ with continuous derivatives up to order $k$; $C^{1,2}([0,T]\times\mathcal{O})$ is the space of functions $u$ on $0,T]\times\mathcal{O}$ for which $\partial_{t}u$, $\partial_{x_{i}}u$, and $\partial_{x_{i}x_{j}}u$ (for $1\leq i,j\leq n$) exist and are continuous. For $u\in C^{2}(\mathcal{O})$, denote by $u_{x}$ its gradient and by $u_{xx}$ its Hessian. Finally, for $a,b\in\mathbb{R}$ let $a\wedge b:=\min\{a,b\}$, and let $\mathbf{1}_{A}(x)$ be the indicator function of a set $A$.
@@ 2.1 Exploratory formulation in reinforcement learning
For a fixed time horizon $T\in(0,\infty)$, let $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ be a filtered probability space with the filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\in[0,T]}$ satisfying the usual conditions. On this probability space, the process $W=(W_{t})_{t\in[0,T]}$ is an $m$-dimensional Brownian motion and the process ${N}=(N(t,z,r);\penalty 10000\ (z,r)\in\mathds{R}^{n}\times\mathds{R}_{+})_{t\in[0,T]}$ is a Poisson point process independent of $W$, having the intensity measure $dt\nu(dz)dr$ for a $\sigma$-finite measure $\nu$ on ${\cal B}(\mathds{R}^{n}\setminus\{0\})$ satisfying $\int_{\mathds{R}^{n}\setminus\{0\}}(|z|^{2}\wedge 1)\nu(dz)<\infty$. Let us consider the controlled jump-diffusion process that $X_{0}^{u}=x\in\mathds{R}^{n}$, and for $t\in(0,T]$,
$dX_{t}^{u}$ $=b(t,X_{t}^{u},u_{t})dt+\sigma(t,X_{t}^{u},u_{t})dW_{t}$
$\quad+\int_{\mathds{R}^{n}\times\mathds{R}_{+}}\varphi(t,X_{t-}^{u},u_{t},z){\bf 1}_{\{r\leq\lambda(t,X_{t-}^{u},u_{t})\}}N(dt,dz,dr),$ (2.1)
where $u=(u_{t})_{t\in[0,T]}$ is an $\mathbb{F}$-predictable process taking values on $U\subseteq\mathds{R}^{d}$, and the set of admissible controls is denoted by $\mathcal{U}$. Here, $b(t,x,u):[0,T]\times\mathds{R}^{n}\times U\to\mathds{R}^{n}$, $\sigma(t,x,u):[0,T]\times\mathds{R}^{n}\times U\to\mathds{R}^{n\times m}$, $\varphi(t,x,u,z):[0,T]\times\mathds{R}^{n}\times U\times\mathds{R}^{n}\to\mathds{R}^{n}$ and $\lambda(t,x,u):[0,T]\times\mathds{R}^{n}\times U\to\mathds{R}_{+}$ are assumed to be measurable. Here, we consider the general jump intensity that is allowed to depend on the state and control.
@@
where $f(t,x,u):[0,T]\times\mathds{R}^{n}\times U\to\mathds{R}$ stands for the running reward function, $g(x):\mathds{R}^{n}\to\mathds{R}$ is the terminal reward function, and the expectation operator $\mathbb{E}_{t}^{\mathbb{P}}[\cdot]:=\mathbb{E}^{\mathbb{P}}\left[\cdot|X_{t}=x\right]$. Given the full knowledge of the coefficients $b,\sigma,\varphi,f,g$, and the intensity measure $\nu$ in (2.1)-(2.2), the classical methods such as the dynamic programming principle and the stochastic maximum principle can be employed to solve the above optimal control problem (2.1)-(2.2). However, in reality, the decision maker may have limited or no information about the environment (i.e., $b,\sigma,\varphi,f,g,\nu$ are unknown). The reinforcement learning approach provides an efficient way to learn the optimal control in (2.1)-(2.2) in the unknown environment through the repeated trial-and-error procedure by taking actions and interacting with the environment. Specifically, he tries a sequence of actions $u=(u_{t})_{t\in[0,T]}$ and observe the corresponding state process $X=(X^{u}_{t})_{t\in[0,T]}$ along with a stream of running rewards $(f(t,X^{u}_{t},u_{t}))_{t\in[0,T]}$ and the terminal reward $g(X_{T}^{u})$, and continuously update and improve his or her actions based on these observations.
@@
To describe the exploration step in reinforcement learning, we can randomize the action $u$ and consider its probability density function. Let ${\cal P}(U)$ be the set of probability density functions on $U$ and $\pi=(\pi_{t})_{t\in[0,T]}$ be a given stochastic feedback policy with $\pi_{t}=\pi(\cdot|t,x)\in{\cal P}(U)$ for any $t\in[0,T]$. For a Borel space $(E,{\cal B}(E))$, consider another probability space $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi})$ and a measurable function $\phi:[0,T]\times\mathbb{R}^{n}\times E\to U$ such that, for all $(t,x)\in[0,T]\times\mathbb{R}^{n}$, the $U$-valued random variable $\omega\to\phi(t,x,\xi(\omega))$ has the probability density $\pi(\cdot|t,x)$. Let $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$ and $(\Omega^{\xi_{n_{0}}},\mathcal{F}^{\xi_{n_{0}}},\mathbb{P}^{\xi_{n_{0}}},\xi_{n_{0}})_{n_{0}\in\mathbb{N}_{0}}$ be independent copies of $(\Omega^{\xi},\mathcal{F}^{\xi},\mathbb{P}^{\xi},\xi)$. Introduce the following probability space with the form given by
@@
$(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{Q}):=\left(\Omega\times\prod_{n_{0}=0}^{\infty}\Omega^{\xi_{n_{0}}},\mathcal{F}\otimes\bigotimes_{n_{0}=0}^{\infty}\mathcal{F}^{\xi_{n_{0}}},\mathbb{P}\otimes\bigotimes_{n_{0}=0}^{\infty}\mathbb{P}^{\xi_{n_{0}}}\right),$ (2.3)
@@
and denotes $\mathbb{F}^{\prime}=(\mathcal{F}_{t}\vee\mathcal{F}^{\prime}_{t})_{t\in[0,T]}$. Furthermore, consider the time grid $G_{t:T}:=\{t=s_{0}<s_{1}<\cdots<s_{n_{0}}=T\}$ and sample actions from $\pi$ only at the grid points in $G_{t:T}$. The state process satisfies, for all $i=0,\ldots,n_{0}-1$ and $s\in[s_{i},s_{i+1}]$,
$X_{s}^{u}$ $=X_{s_{i}}^{u}+\int_{s_{i}}^{s}b(\ell,X_{\ell}^{u},u_{s_{i}})d\ell+\int_{s_{i}}^{s}\sigma(\ell,X_{\ell}^{u},u_{s_{i}})dW_{\ell}$
$\qquad+\int_{s_{i}}^{s}\int_{\mathds{R}^{n}\times\mathds{R}_{+}}\varphi(\ell,X_{\ell-}^{u},u_{s_{i}},z){\bf 1}_{\{r\leq\lambda(\ell,X_{\ell-}^{u},u_{s_{i}})\}}N(d\ell,dz,dr),$ (2.4)
$u_{s_{i}}$ $=\phi(s_{i},X_{s_{i}},\xi_{i}).$
@@
$dX_{s}^{\pi}$ $=b(s,X_{s}^{\pi},u_{s}^{\pi})ds+\sigma(s,X_{s}^{\pi},u_{s}^{\pi})dW_{s}+\int_{\mathds{R}^{n}\times\mathds{R}_{+}}\varphi(s,X_{s-}^{\pi},u_{s}^{\pi},z){\bf 1}_{\{r\leq\lambda(s,X_{s-}^{u},u_{s})\}}N(ds,dz,dr),$
$u_{s}^{\pi}$ $=u_{\delta(s)}\penalty 10000\ \text{with\penalty 10000\ $\delta(s)=s_{i}$ when $s\in[s_{i},s_{i+1})$}.$ (2.5)
To encourage the exploration in RL, we consider the so-called Tsallis entropy regularization. Then, the objective functional is defined by
@@
$l_{p}(z)=\begin{cases}\displaystyle\frac{1}{p-1}(1-z^{p-1}),&p>1,\\[6.00006pt] \displaystyle\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ -\ln z,&p=1.\end{cases}$ (2.7)
@@
However, the representation (2.1)-(2.6) cannot be applied to derive exploratory HJB equation directly from the point of view of DPP. To this purpose, it is necessary to provide the relaxed version of the control problem. Inspired by Kushner (2000) and Gao et al. (2024) , let us introduce the process $\tilde{X}=(\tilde{X}_{s})_{s\in[t,T]}$ satisfying the following SDE, $\tilde{X}_{t}^{\pi}=x$, and for $s\in(t,T]$,
@@
$\quad+\int_{U}\int_{\mathds{R}^{n}\times\mathds{R}_{+}}\varphi(s,\tilde{X}_{s-}^{\pi},u,z){\bf 1}_{\{r\leq\lambda(s,\tilde{X}_{s-}^{\pi},u)\}}\mathcal{N}(ds,dz,dr,du),$ (2.8)
where $B=(B_{t})_{t\in[0,T]}$ is a scalar Brownian motion and $\mathcal{N}(ds,dz,dr,du)$ is a Poisson random measure on $\mathds{R}_{+}\times{\mathds{R}}^{n}\times\mathds{R}_{+}\times U$ with compensator $ds\nu(dz)dr\pi_{s}(u)du$ independent of $B$. An interesting finding is that, for the pure jump controlled state model, the representation (2.1) of the relaxed controlled state process can be applied to derive exploratory HJB equations directly from the point of view of DPP, which is different from the controlled diffusion case as in Wang et al. (2020). Therefore, we can formulate our reinforcement learning problem for the jump-diffusion controlled model (2.1) based on the relaxed control form (2.1). Thus, our reinforcement learning problem associated with the jump-diffusion controlled state process (2.1) can be stated as follows:
@@
$\penalty 10000\ \penalty 10000\ {\rm s.t.}\penalty 10000\ \tilde{X}_{s}^{\pi}=x+\int_{t}^{s}\tilde{b}(\ell,\tilde{X}_{\ell}^{\pi},\pi_{\ell})d\ell+\int_{t}^{s}\tilde{\sigma}(\ell,\tilde{X}_{\ell}^{\pi},\pi_{\ell})dB_{\ell}$
$\qquad\qquad\qquad+\int_{t}^{s}\int_{U}\int_{\mathds{R}^{n}\times\mathds{R}_{+}}\varphi(\ell,\tilde{X}_{\ell-}^{\pi},u,z){\bf 1}_{\{r\leq\lambda(\ell,\tilde{X}_{\ell-}^{\pi},u)\}}\mathcal{N}(d\ell,dz,dr,du).$
Here, $\Pi_{t}$ is the set of admissible (randomized) policies on $U$ and the coefficients $\tilde{b},\tilde{\sigma},\tilde{f},\tilde{l}_{p}$ are defined by, for $(s,x,\pi)\in[0,T]\times\mathds{R}^{n}\times{\cal P}(U)$,
$\tilde{b}(s,x;\pi):=\int_{U}b(s,x,u)\pi(u|s,x)du,\quad\tilde{\sigma}(s,x;\pi):=\sqrt{\int_{U}\sigma^{2}(s,x,u)\pi(u|s,x)du},$
$\tilde{f}(s,x;\pi):=\int_{U}f(s,x,u)\pi(u|s,x)du,\quad\tilde{l}_{p}(s,x;\pi):=\int_{U}l_{p}(\pi(u|s,x))\pi(u|s,x)du.$
@@
    $\pi(\cdot|s,x)\in{\cal P}(U)$ for all $(s,x)\in[t,T]\times\mathds{R}^{n}$, and $\pi(\cdot|\cdot,\cdot):U\times[t,T]\times\mathds{R}^{n}\to\mathds{R}$ is a measurable function;
@@
    $\int_{U}\left|\pi(u|s,x_{1})-\pi\left(u|s,x_{2}\right)\right|du\leq C\left|x_{1}-x_{2}\right|,\quad\forall x_{1},x_{2}\in\mathds{R}^{n}.$